OpenStudy (anonymous):

The intensity of an x-ray beam at any distance from the source varies inversely as the square of that distance. If the intensity of an x-ray beam is 30 R/hour when the object is 2 inches from the source, determine the level of intensity when an object is 5 inches from the source. (Note: R/hour stands for Roentgens/hour. A Roentgen is a unit for measuring the amount of x-ray or gamma radiation in the air.)

OpenStudy (anonymous):

it varies inversely so I know you divide I think 30 by 15?

OpenStudy (anonymous):

k= 60?????????

OpenStudy (anonymous):

Let \(i\) represent the intensity and let \(\delta\) represent the distance. So we have: \[i\propto\frac{1}{\delta^2}\] That is to say: \[i=\frac{\chi}{\delta^2}\] Where \(\chi\) is a constant. So we know: \[30=\frac{\chi}{2^2}\to30=\frac{\chi}{4}\to\chi=30(4)\to\chi=120\] We also know that: \[i=\frac{\chi}{5}\to i=\frac{120}{5}=24R/hr\]

OpenStudy (anonymous):

that's what I thought too but here are my choices: A. 15/ 2 B. 15 C. 60 D. 120

OpenStudy (anonymous):

OH SHOOT! haha i made a mistake: we have, \[i=\frac{\chi}{5^2}=\frac{\chi}{25}=\frac{120}{25}=\frac{24}{5}=4.8R/hr\]

OpenStudy (anonymous):

does that fall into any of the choices?

OpenStudy (anonymous):

Nope...something doesn't make sense.